# Harnessing Vision-Language Models for Time Series Anomaly Detection

### v1

@@ Appendix B Additional Experimental Results
B.1 Full Ablation Analysis
Table 5 presents the full ablation results, extending the analyses from Section 4.3 with two additional experiments. First, we evaluate an all-pairs cross-patch comparison—computing the minimum cosine dissimilarity between every patch in the test window and every patch across all other windows (instead of using a single median reference). As expected, all-pairs matching yields higher F1-max across multiple threshold quantiles, though at the expense of an increased memory usage (as we need to store embeddings for all windows). The default median-reference approach (reported in the main paper) therefore remains preferable for its efficiency.
Second, we vary the quantile $q$ used to collapse the 2-D anomaly map to 1-D (Section 3.1). In addition to the standard $q=25\%$, we test $q=12.5\%$ and $q=50\%$. Reducing $q$ to 12.5% produces nearly identical F1-max, while increasing $q$ to 50% improves detection on signals with extended, high-amplitude anomalies (e.g. long spikes). All other variants—patch-level embedding, cross-patch matching ablations, multi-scale embedding, and omission of the ViT4TS screening stage—reproduce the trends reported in Section 4.3, confirming that (i) fine-grained patch features, (ii) flexible cross-patch comparision, (iii) multi-scale feature extraction are each crucial to our performance gains.
@@ B.1 Full Ablation Analysis
VLM4TS Prompt You are an expert in both time-series analysis and multimodal (vision + language) reasoning. You will be shown: 1. A plot of raw time-series data X-axis: time step index Y-axis: signal value over time 2. Preliminary “vision-based” anomaly windows A list of intervals detected by a coarse, purely visual model (may include false positives and false negatives) Your goal is to integrate both sources—the visual plot and the preliminary windows—and produce a refined, final anomaly detection for the entire series. Specifically: Eliminate any preliminary windows that look anomalous in isolation but are consistent with the overall trend. Add any intervals that the visual model missed but which break temporal continuity or exhibit clear statistical irregularities (spikes, level shifts, abrupt changes). Response format Reply only with a JSON object containing these fields:
@@
where: "interval_index": an array of [start, end] pairs (inclusive indices). "confidence": a parallel array of integers (1–3 scale). "abnormal_description": a single paragraph (less than 100 words) summarizing why these intervals are anomalous. Confidence scale: 1 = Low confidence: ambiguous or very subtle deviation. 2 = Medium confidence: clear local irregularity but moderate global uncertainty. 3 = High confidence: strong statistical or contextual evidence of anomaly. Important: Estimate interval boundaries using the tick marks on the x-axis as precisely as possible. The very first segment may appear atypical due to slicing; do not flag it without clear anomaly evidence. Do not include any extra keys or commentary—only the JSON object above.
B.2 VLM Comparison
Table 6 and 7 report the maximum F1 scores of VLM4TS when using different VLM backbones for the verification stage. Overall, GPT-4o achieves the highest mean F1, demonstrating balanced performance across domains. Claude-3.5 performs competitively with GPT-4o, particularly on AWS, Traffic and Art datasets. Gemini-2.0 shows its strength on the Yahoo A3 and A4 benchmarks as well as the NAB Traffic subset. We note that Gemini-2.0 is more aggressive when selecting anomalies thus it boosts recall on high-frequency anomalies like those in A3 and A4. These variations show that each VLM’s internal verification heuristics—whether conservative or aggressive—can shift the precision–recall trade-off depending on the anomaly characteristics and domain.
Table 6: Max-F1 scores of VLM4TS with different VLMs on NASA and YAHOO dataset groups, with percentage change relative to GPT-4o.
NASA YAHOO
Method MSL SMAP A1 A2 A3 A4
GPT-4o 0.619 0.773 0.733 0.901 0.497 0.474
Claude-3.5 0.588 ($-5.0\%$) 0.723 ($-6.5\%$) 0.708 ($-3.4\%$) 0.887 ($-1.6\%$) 0.530 ($+6.6\%$) 0.490 ($+3.4\%$)
Gemini-2.0 0.554 ($-10.4\%$) 0.754 ($-2.5\%$) 0.723 ($-1.4\%$) 0.865 ($-4.0\%$) 0.531 ($+6.8\%$) 0.496 ($+4.6\%$)
Table 7: Max-F1 scores of VLM4TS with different VLMs on NAB datasest group, with percentage change relative to GPT-4o.
NAB
Method AdEx Tweets Traffic AWS Art
GPT-4o 0.727 0.686 0.632 0.488 0.714
Claude-3.5 0.625 ($-14.1\%$) 0.617 ($-10.0\%$) 0.600 ($-4.9\%$) 0.543 ($+11.2\%$) 0.674 ($-5.6\%$)
Gemini-2.0 0.583 ($-19.8\%$) 0.643 ($-6.3\%$) 0.703 ($+11.2\%$) 0.439 ($-10.1\%$) 0.625 ($-12.4\%$)
B.3 Elapsed Time Comparison
@@
In Table 8, we compare end-to-end elapsed time (measured on an NVIDIA V100 GPU) for ViT4TS, VLM4TS against both scratch-trained and time-series-pretrained TSAD baselines. We note that different class of model follows its different hyperparameter settings based on the protocol [45]: forecasting methods (ARIMA, ATrans, TimesFM, LSTM-DT, UniTS) use window size of 250 and step size of 1, and reconstruction methods (AER, TadGAN, VAE, LSTM-AE) use window size of 100 and step size of 1; ViT4TS employs a 224×224 window with a step size of 56 (one-quarter window). The reason we choose such a step size is that for each image we are applying multi-scale feature extraction and patch-level cross-patch comparison, which allows for larger step size. For scratch-trained networks, timings include both training and inference; for pretrained models (UniTS, TimesFM) and our methods, we report inference time only. For comparison in elapsed time with language model-based methods, see Table 2. The Token usage and computation time is adjusted based on the window length and stride size, making sure that in Table 2 all the metrics are directly comparable. Despite its multi-scale screening and VLM verification stages, ViT4TS runs in seconds on moderate-length series and VLM4TS incurs only a few seconds per candidate via the VLM API, indicating that our two-stage pipeline is sufficiently efficient for deployment in real-world industrial monitoring systems.
@@
Table 8: Elapsed time comparison of VLM4TS versus trained-from-scratch and time-series-pretrained baselines on benchmark datasets. Different methods follow different hyperparameter settings, with details discussed in Appendix B.3.
@@
B.4 Recall and Precision Comparison between ViT4TS and VLM4TS
Tables 9, 10, and 11 present precision and recall for ViT4TS and VLM4TS across all 11 benchmarks under three threshold settings. As expected, the verification stage usually increases precision: in most datasets, VLM4TS reduces false positives generated by the high-recall ViT4TS screening, particularly at lower thresholds where ViT4TS produces many candidate intervals. This precision gain comes with a modest recall reduction, since VLM4TS filters out some true positives; however, the net effect is an improved F1-max in most datasets. The exceptions are A3 and A4 datasets, where the abundance of anomalies makes conservative filtering less beneficial.
On the other hand, VLM4TS also boosts recall in datasets dominated by contextual anomalies—like MSL, SMAP, Art—by flagging additional intervals that ViT4TS missed due to its limited local window. In these cases, the model’s global reasoning corrects false negatives and yields a more balanced precision–recall profile. Overall, VLM4TS improves the screening baseline’s F1-max in nine out of eleven tasks, showing the necessity of the two-stage localization and verification design.
B.5 Embedding Analysis
@@
B.6 Full F1 scores for Table 1
@@
Time-series-pretrained TimesFM 0.410 0.500 0.564 0.434 0.686 0.602 0.127 0.182 0.198
UniTS 0.531 0.550 0.561 0.560 0.715 0.723 0.130 0.167 0.120
@@
Time-series-Pretrained TimesFM 0.224 0.234 0.000 0.243 0.194 0.202 0.152 0.270 0.400 0.467 0.333 0.343
UniTS 0.182 0.000 0.000 0.228 0.246 0.214 0.246 0.326 0.312 0.479 0.383 0.323
@@
Time-series-pretrained TimesFM 0.221 0.516 0.554 0.215 0.694 0.621 0.120 0.044 0.015 0.107 0.066 0.019
UniTS 0.248 0.520 0.605 0.205 0.760 0.748 0.121 0.126 0.107 0.110 0.090 0.064

### v2

@@ Appendix B Additional Experimental Results
Full Ablation Analysis
Table 5 presents the full ablation results, extending the analyses from Section 4.3 with two additional experiments. First, we evaluate an all‐pairs cross‐patch comparison—computing the minimum cosine dissimilarity between every patch in the test window and every patch across all other windows (instead of using a single median reference). As expected, all‐pairs matching yields higher F1‐max across multiple threshold quantiles, though at the expense of an increased memory usage (as we need to store embeddings for all windows). The default median‐reference approach (reported in the main paper) therefore remains preferable for its efficiency.
Second, we vary the quantile $q$ used to collapse the 2-D anomaly map to 1-D (Section 3.1). In addition to the standard $q=25\%$, we test $q=12.5\%$ and $q=50\%$. Reducing $q$ to 12.5% produces nearly identical F1‐max, while increasing $q$ to 50% improves detection on signals with extended, high‐amplitude anomalies (e.g. long spikes). All other variants—patch‐level embedding, cross‐patch matching ablations, multi‐scale embedding, and omission of the ViT4TS screening stage—reproduce the trends reported in Section 4.3, confirming that (i) fine‐grained patch features, (ii) flexible cross-patch comparision, (iii) multi‐scale feature extraction are each crucial to our performance gains.
@@
Figure 6: Multivariate extension of our framework. Left: ViT4TS detects candidate anomalies independently for each channel. Bottom: VLM4TS receives both the multivariate visualization and a textual summary of all channel–wise candidates, enabling cross–channel visual understanding for a refined detection. Right: final multivariate anomaly intervals produced by VLM4TS.
Multivariate Extension.
While our main experiments focus on univariate time series to highlight the key idea of the two–stage framework, the method extends naturally to multivariate settings. The key idea is to keep the first stage channel–specific and allow the second stage to perform cross–channel visual understanding and reasoning to produce the final detection.
Figure 6 illustrates this procedure. On the left, ViT4TS is applied independently to each channel, producing channel–wise candidate intervals based on local shape deviations. These raw candidates often include false positives or isolated fluctuations. In the second stage, VLM4TS receives (i) an aggregated textual summary of all channel–wise candidates and (ii) a stacked‐subplot visualization of the full multivariate series. As shown in the bottom panel in Figure 6, with access to this global view, the VLM is able to reason across channels, retaining intervals that exhibit coordinated cross–channel anomalies while discarding channel–specific noise. The right panel shows the resulting multivariate anomaly intervals.
We conduct preliminary experiments on the multivariate MSL benchmark using this pipeline. Incorporating cross–channel context raises the Max-F1 from 0.619 (univariate) to 0.660, suggesting that VLM–based integration across channels provides additional benefit beyond single–channel counterpart. This multivariate extension preserves the design of the univariate framework while leveraging the VLM’s ability to fuse information from different time series channels.
VLM4TS Prompt You are an expert in both time‐series analysis and multimodal (vision + language) reasoning. You will be shown: 1. A plot of raw time‐series data X‐axis: time step index Y‐axis: signal value over time 2. Preliminary “vision‐based” anomaly windows A list of intervals detected by a coarse, purely visual model (may include false positives and false negatives) Your goal is to integrate both sources—the visual plot and the preliminary windows—and produce a refined, final anomaly detection for the entire series. Specifically: Eliminate any preliminary windows that look anomalous in isolation but are consistent with the overall trend. Add any intervals that the visual model missed but which break temporal continuity or exhibit clear statistical irregularities (spikes, level shifts, abrupt changes). Response format Reply only with a JSON object containing these fields:
@@
where: "interval_index": an array of [start, end] pairs (inclusive indices). "confidence": a parallel array of integers (1–3 scale). "abnormal_description": a single paragraph (less than 100 words) summarizing why these intervals are anomalous. Confidence scale: 1 = Low confidence: ambiguous or very subtle deviation. 2 = Medium confidence: clear local irregularity but moderate global uncertainty. 3 = High confidence: strong statistical or contextual evidence of anomaly. Important: Estimate interval boundaries using the tick marks on the x‐axis as precisely as possible. The very first segment may appear atypical due to slicing; do not flag it without clear anomaly evidence. Do not include any extra keys or commentary—only the JSON object above.
Elapsed Time Comparison
@@
In Table 6, we compare end-to-end elapsed time (measured on an NVIDIA V100 GPU) for ViT4TS, VLM4TS against both scratch-trained and time-series-pretrained TSAD baselines. We note that different class of model follows its different hyperparameter settings based on the protocol (alnegheimish2024orionbench): forecasting methods (ARIMA, ATrans, TimesFM, LSTM-DT, UniTS) use window size of 250 and step size of 1, and reconstruction methods (AER, TadGAN, VAE, LSTM-AE) use window size of 100 and step size of 1; ViT4TS employs a 224×224 window with a step size of 56 (one-quarter window). The reason we choose such a step size is that for each image we are applying multi-scale feature extraction and patch-level cross‐patch comparison, which allows for larger step size. For scratch-trained networks, timings include both training and inference; for pretrained models (UniTS, TimesFM) and our methods, we report inference time only. For comparison in elapsed time with language model-based methods, see Table 2. The Token usage and computation time is adjusted based on the window length and stride size, making sure that in Table 2 all the metrics are directly comparable. Despite its multi‐scale screening and VLM verification stages, ViT4TS runs in seconds on moderate‐length series and VLM4TS incurs only a few seconds per candidate via the VLM API, indicating that our two-stage pipeline is sufficiently efficient for deployment in real-world industrial monitoring systems.
@@
Table 6: Elapsed time comparison of VLM4TS versus trained-from-scratch and time-series-pretrained baselines on benchmark datasets. Different methods follow different hyperparameter settings, with details discussed in the Appendix.
@@
Embedding Analysis
@@
Recall and Precision Comparison between ViT4TS and VLM4TS
Tables 7, 8, and 9 present precision and recall for ViT4TS and VLM4TS across all 11 benchmarks under three threshold settings. As expected, the verification stage usually increases precision: in most datasets, VLM4TS reduces false positives generated by the high‐recall ViT4TS screening, particularly at lower thresholds where ViT4TS produces many candidate intervals. This precision gain comes with a modest recall reduction, since VLM4TS filters out some true positives; however, the net effect is an improved F1-max in most datasets. The exceptions are A3 and A4 datasets, where the abundance of anomalies makes conservative filtering less beneficial.
On the other hand, VLM4TS also boosts recall in datasets dominated by contextual anomalies, like MSL, SMAP, Art, by flagging additional intervals that ViT4TS missed due to its limited local window. In these cases, the model’s global reasoning corrects false negatives and yields a more balanced precision–recall profile. Overall, VLM4TS improves the screening baseline’s F1-max in nine out of eleven tasks, showing the necessity of the two‐stage localization and verification design.
@@
Choice of Patch Size and Its Effect on Screening Performance
In Table 1 (Max-F1) and Table 9 (precision–recall), the ViT4TS screening stage shows noticeably lower recall on the A3 and A4 datasets, resulting in reduced Max-F1 compared with the best TSAD baselines. To understand this behavior, we analyze how ViT4TS responds to different anomaly structures. As shown in Figure 8, anomalies in A3 and A4 are short, dense, and concentrated within narrow intervals. Under the default patch configuration (window size 224), several such point anomalies fall within a single window, producing coarse anomaly maps that under-represent the fine-scale deviations.
To further understand the effect of patch size, we evaluated a smaller patch size (window size 28), which offers substantially finer temporal granularity for cross-patch comparison. This adjustment yields a marked improvement, increasing Max-F1 from 0.614 to 0.906 on A3 and from 0.565 to 0.844 on A4, surpassing all baselines reported in Table 1. Figure 8 also illustrates this effect. Using the default patch size (window size 224; middle panel), anomaly scores merge multiple point anomalies into broad peaks, yielding coarse localization and reduced sensitivity. In contrast, the smaller patch configuration (window size 28; bottom panel) produces sharply localized anomaly responses that align closely with the rapid, fine-scale deviations in the signal. These results underscore the importance of patch and window size selection in capturing subtle, high-frequency anomaly patterns. Developing data-driven or adaptive strategies for patch sizing is therefore a promising direction for further improving the ViT4TS screening stage.
Figure 8: Patch-size effect on ViT4TS for anomaly scoring on A3 TS2. Top: original time series with ground-truth anomalies (red). Middle: anomaly scores produced using the default patch size 224, which smooths over short and densely occurring anomalies. Bottom: anomaly scores using a smaller patch size 28, yielding substantially improved localization.
Choice of VLM Backbone
Table 10 and 11 report the Max-F1 scores of VLM4TS when using with different VLM backbones, including the close-sourced Gemini-2.0 and Claude-3.5 as well as the open-sourced Qwen-VL-2.5–72B. While these models exhibit different verification behaviors across tasks, the overall performance remains at or above state-of-the-art TSAD baselines presented in Table 1.
Table 10: Max‐F1 scores of VLM4TS with different VLMs on NASA and YAHOO dataset groups, with percentage change relative to GPT-4o.
NASA YAHOO
Method MSL SMAP A1 A2 A3 A4
GPT-4o 0.619 0.773 0.733 0.901 0.497 0.474
Claude-3.5 0.588 ($5.0\%$) 0.723 ($6.5\%$) 0.708 ($3.4\%$) 0.887 ($1.6\%$) 0.530 ($6.6\%$) 0.490 ($3.4\%$)
Gemini-2.0 0.554 ($10.4\%$) 0.754 ($2.5\%$) 0.723 ($1.4\%$) 0.865 ($4.0\%$) 0.531 ($6.8\%$) 0.496 ($4.6\%$)
Qwen-2.5-72B 0.606 ($2.1\%$) 0.805 ($4.1\%$) 0.747 ($1.9\%$) 0.919 ($2.0\%$) 0.492 ($1.0\%$) 0.531 ($12.0\%$)
Table 11: Max‐F1 scores of VLM4TS with different VLMs on NAB datasest group, with percentage change relative to GPT-4o.
NAB
Method AdEx Tweets Traffic AWS Art
GPT-4o 0.727 0.686 0.632 0.488 0.714
Claude-3.5 0.625 ($14.1\%$) 0.617 ($10.0\%$) 0.600 ($4.9\%$) 0.543 ($11.2\%$) 0.674 ($5.6\%$)
Gemini-2.0 0.583 ($19.8\%$) 0.643 ($6.3\%$) 0.703 ($11.2\%$) 0.439 ($10.1\%$) 0.625 ($12.4\%$)
Qwen-2.5-72B 0.667 ($8.3\%$) 0.630 ($8.2\%$) 0.649 ($2.7\%$) 0.500 ($2.5\%$) 0.857 ($20.0\%$)
Comparison with Spectrogram-Based TSAD Approaches.
Spectrograms provide an alternative 2D representation of time series by encoding time–frequency structure, offering some global context. However, their use in zero-shot TSAD is restricted. Spectrogram characteristics are highly domain dependent and considerably less interpretable than line plots. Second, their visual appearance differs substantially from natural images used in large-scale vision pretraining, which weakens the transferability of off-the-shelf vision encoders and VLMs.
It’s also worth noting that spectrograms do not resolve the resolution–context dilemma: long sequences must still be divided into fixed windows before being processed, as in ITF-TAD (namura2024training), leaving the trade-off between fine local resolution and global temporal coverage unresolved. Our two-stage design directly addresses this challenge by decoupling the tasks: ViT4TS performs fine-grained localization on high-resolution line plots, while VLM4TS conducts global verification with access to the full temporal horizon.
We compare ViT4TS and VLM4TS against the spectrogram-based ITF-TAD baseline. As shown in Table 12, ITF-TAD underperforms ViT4TS across most datasets, suggesting that spectrogram statistics are not well aligned with vision models pretrained on natural images. Furthermore, ITF-TAD falls far short of VLM4TS as it lacks a mechanism to combine fine-grained localization with global temporal understanding.
Table 12: Comparison of ITF-TAD (spectrogram) versus ViT4TS (line-plot) and VLM4TS.
Method NAB NASA YAHOO
Art AWS AdEx Traf Tweets MSL SMAP A1 A2 A3 A4 $\boldsymbol{\mu\pm\sigma}$
[…]
Full F1 scores for Table 1
@@
Time-series-pretrained UniTS 0.531 0.550 0.561 0.560 0.715 0.723 0.130 0.167 0.120
TimesFM 0.410 0.500 0.564 0.434 0.686 0.602 0.127 0.182 0.198
TimesFM2 0.391 0.512 0.556 0.202 0.608 0.676 0.158 0.156 0.171
@@
Time-series-Pretrained UniTS 0.182 0.000 0.000 0.228 0.246 0.214 0.246 0.326 0.312 0.479 0.383 0.323
TimesFM 0.224 0.234 0.000 0.243 0.194 0.202 0.152 0.270 0.400 0.467 0.333 0.343
TimesFM2 0.255 0.235 0.000 0.233 0.177 0.176 0.196 0.286 0.364 0.386 0.381 0.207
@@
Time-series-pretrained UniTS 0.248 0.520 0.605 0.205 0.760 0.748 0.121 0.126 0.107 0.110 0.090 0.064
TimesFM 0.221 0.516 0.554 0.215 0.694 0.621 0.120 0.044 0.015 0.107 0.066 0.019
TimesFM2 0.271 0.577 0.594 0.313 0.692 0.630 0.202 0.084 0.013 0.181 0.106 0.032